OpenStudy (anonymous):

The product of two consecutive positive odd integers equals 91 more than eight times the first integer. Find both integers.

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

i cna do it

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

15 and 17

OpenStudy (anonymous):

let first integer is X so other is X+2. X(X+2)=8X+91 X^2-6X-91=0 (X-13)(X-7)=0 X=13, and x=7 x+2=15 and x+2=9 numbers are 13 and 15 0r 7 and 9

OpenStudy (anonymous):

sorry -7 and 5